A nuclear power plant operates at 65% of its maximum theoretical (Carnot)

efficiency between temperatures of 660 0 C and 330 0 C. If the plant produces electric

energy at a rate of 1.2 GW, how much exhaust heat is discharged per hour?

[Ans = 1.4 X 10 13 J/h]


1
Expert's answer
2022-01-19T05:59:57-0500

"\dot Q_2=\gamma W_1\frac{T_2}{T_1}=1.4\cdot 10^{13}~\frac Jh."